### v1

# Revealing the $\chi_{\rm eff}$-$q$ correlation among Coalescing Binary Black Holes and Evidence for AGN-driven Hierarchical Mergers

Abstract
The origin of the correlation between the effective spins ($\chi_{\rm eff}$) and mass ratios ($q$) of LIGO-Virgo-KAGRA’s binary black holes (BBHs) is still an open question. Motivated by recent identification of two subpopulations of the BBHs, in this work we investigate the potential $\chi_{\rm eff}-q$ correlation for each subpopulation. Surprisingly, the $\chi_{\rm eff}$-$q$ correlation vanishes for the low-mass subpopulation if we introduce a second $\chi_{\rm eff}$ distribution for the high-mass subpopulation likely originating from hierarchical mergers. The first subpopulation has a narrow $\chi_{\rm eff}$ distribution peaking at $\sim 0.05$, whose primary-mass function cuts off at $\sim 45M_{\odot}$, in agreement with first-generation BBHs. The second $\chi_{\rm eff}$ distribution is broad and peaks at $\mu_{\chi,2}=0.35^{+0.18}_{-0.22}$, consistent with the expectation of hierarchical mergers formed in the disks of active galactic nucleus (AGNs). We infer $\mu_{\chi,2}>0$ at 98.7% credible level, and a symmetric $\chi_{\rm eff}$ distribution for the second subpopulation is disfavored by $\mathcal{B}\sim 5$. However, negative values of $\chi_{\rm eff}$ are also measured, indicating that the hierarchical mergers may take place within both star clusters and AGN disks. We find a Bayes factor of $\ln\mathcal{B}=5.2$ for two distinct $\chi_{\rm eff}$ distributions relative to single $\chi_{\rm eff}$ distribution that conditioned on mass ratios. Therefore we conclude that the $\chi_{\rm eff}$-$q$ correlation in the entire population can be explained as the superposition of two subpopulations. Additionally, we suggest to use a flexible mass function to reduce the bias in $\chi_{\rm eff}$-$q$ correlation that may be introduced by model mis-specification.
Binary Black Holes; Gravitational Waves; Stellar Evolution; Active Galactic Nuclei
software: Bilby (Ashton et al., 2019, version 1.1.4, ascl:1901.011, https://git.ligo.org/lscsoft/bilby/), PyMultiNest (Buchner, 2016, version 2.11, ascl:1606.005, https://github.com/JohannesBuchner/PyMultiNest).
{CJK*}
UTF8gbsn
1 Introduction
The coalescing binary black holes (BBHs) will provide clues about their formation and evolutionary processes through the parameters of these systems (Mandel & Farmer, 2022). In addition to characterizing the marginalized distributions of these parameters, it is also important to investigate the correlations among them (Heinzel et al., 2024; Callister, 2024), including mass versus spin (e.g. Tiwari & Fairhurst, 2021; Wang et al., 2022; Li et al., 2024b, a), mass versus mass ratio (Li et al., 2022), spin versus mass ratio (Callister et al., 2021), redshift versus spin (Biscoveanu et al., 2022), and redshift versus mass versus spin (Tiwari, 2022; Guo et al., 2024).
Callister et al. (2021), for the first time, reported the anti-correlation between the effective spins ($\chi_{\rm eff}$) and mass ratios ($q$) of BBHs with data from GWTC-2 (Abbott et al., 2019, 2021a). This anti-correlation was subsequently confirmed by Abbott et al. (2023a) with data from GWTC-3 (Abbott et al., 2019, 2021a, 2024, 2023b). More detailed analysis (e.g., Adamcewicz & Thrane, 2022; Adamcewicz et al., 2023) had shown that the $\chi_{\rm eff}$-$q$ anti-correlation is not artificial. However, the origin of this anti-correlation is still in debate (Callister et al., 2021; Abbott et al., 2023a).
Many simulations suggested that hierarchical mergers in the disks of active galactic nucleus (AGNs) could explain the observed $\chi_{\rm eff}$-$q$ correlation (e.g., McKernan et al., 2022; Santini et al., 2023; Cook et al., 2024). The spin orientations and orbital angular momenta of binary black holes (BBHs) will be modulated by the disks, causing the mergers to favor positive effective spins. Additionally, the migration traps in AGN disks will produce mergers involving multiple generations of black holes (BHs) that have unequal masses and larger (positive) effective spins. Previously, we found that the coalescing BHs can be divided into two subpopulations with significantly different spin-magnitude versus component-mass distributions, which are nicely consistent with first- and higher-generation BHs (Li et al., 2024b). We also found that a fraction of the the hierarchical mergers have aligned spin tilts and asymmetric mass ratios, which may give rise to the $\chi_{\rm eff}$-$q$ anti-correlation.
However, other simulations suggested that certain formation channels of isolated binaries can produce highly spinning and unequal-mass BBHs, particularly the stable mass transfer formation scenario (Banerjee & Olejak, 2024). Because the progenitors of unequal BBH systems in the stable mass transfer formation scenario are more likely to efficiently shrink their orbits during the second Roche-lobe overflow, which makes them easier to enter the tidal spin-up regime and later merge due to GW emission (Olejak et al., 2024). Therefore, it is possible that there may be a correlation between the effective spin and mass ratio in first-generation (or low-spin) BBHs, or among potential isolated evolution channels (Wang et al., 2022; Godfrey et al., 2023; Li et al., 2024a).
Therefore, in this study, we delve into the question of whether the anti-correlation between $\chi_{\rm eff}$-$q$ in BBHs originates from a superposition of various formation channels / subpopulations (Li et al., 2024b), if it emerges from the evolutionary processes of a single population with a common formation channel (e.g., Banerjee & Olejak, 2024), or alternatively, both effects have contributed to the observed correlation. Additionally, we investigate the $\chi_{\rm eff}$ distribution in the high-mass range that corresponds to the hierarchical mergers, in order to find out whether these events originate from the AGN disks, or alternatively from the star clusters which exhibits a symmetric $\chi_{\rm eff}$ distribution (Payne et al., 2024; Antonini et al., 2024).
This letter is organized as follows: In Section 2 and Section 3, we introduce the methods and the results. In Section 4, we present the conclusions and engage in discussions.
2 Methods
We use hierarchical Bayesian inference to measure the hyperparameters of population model; see Appendix A for details. Following Abbott et al. (2023a), we adopted 69 BBH events with false alarm rates (FAR) $<1{\rm yr}^{-1}$ in GWTC-3 for analysis. The posterior samples for each BBH event are obtained from events-zenodo, and the ‘C01:Mixed’ samples are adopted.
We first use the same $\chi_{\rm eff}-q$ distribution model as that used in Callister et al. (2021) to fit data of GWTC-3 for comparison. Specifically, the mean and width of the $\chi_{\rm eff}$ distribution change linearly with $q$ (referred to as the Base model); see Appendix B.2 for the detailed formula. The rate evolution model is the MD model (Madau & Dickinson, 2014), as defined in the Appendix B.3. To reduce the bias that may be introduced by the mis-modeling of primary-mass distribution, we use the concise non-/semi-parametric formula PowerLaw+Spline (PS; Edelman et al., 2022; Abbott et al., 2023a) instead of the popular parametric formula PowerLaw+Peak (PP; Talbot & Thrane, 2018; Abbott et al., 2021b), see Appendix B.1 for the details of mass functions. We find the PP model is disfavored compared to the PS model by $\ln\mathcal{B}=6$, due to its failure in modeling the additional structures in the primary-mass distribution (Li et al., 2021; Tiwari, 2022; Edelman et al., 2023; Abbott et al., 2023a; Callister & Farr, 2024). With the PP model, we find an anti-correlation between $\chi_{\rm eff}$ and $q$ distribution (i.e., $a<0$) at 99.2% credible level. However, the credibility decreases to 89.8% with the more flexible PS model, indicating that mis-modeling in the primary-mass distribution may introduce bias in measuring the $\chi_{\rm eff}$-$q$ correlation; however, the overall tendency remains unchanged, as shown in Figure 1.
Inspired by the investigation of spin-magnitude versus component-mass distribution (Li et al., 2024b), we introduce another $\chi_{\rm eff}$-distribution to capture the secondary subpopulation of BBHs that is consistent with hierarchical mergers. Two kind of models are applied, one is the Mixture model on primary-mass versus effective-spin distribution,
$\pi_{\rm mix}$ $(m_{1},m_{2},\chi_{\rm eff}|{\bf\Lambda})=P(m_{2}|m_{1};{\bf\Lambda})$ (1)
$[\pi_{1}(m_{1},\chi_{\rm eff}|{\bf\Lambda})(1-r_{2})+\pi_{2}(m_{1},\chi_{\rm eff}|{\bf\Lambda})r_{2}],$
with
$\pi_{1}$ $(m_{1},\chi_{\rm eff}|{\bf\Lambda})=\mathcal{PS}(m_{1}|{\bf\Lambda})P(\chi_{\rm eff}|q;\mu_{\chi,0},\sigma_{\chi,0},a,b)$ (2)
and
$\pi_{2}$ $(m_{1},\chi_{\rm eff}|{\bf\Lambda})=\mathcal{PL}(m_{1}|-\alpha_{2},m_{\rm min,2},m_{\rm max,2})$ (3)
$\mathcal{G}_{[-1,1]}(\chi_{\rm eff}|\mu_{\chi,2},\sigma_{\chi,2})$
where $r_{2}$ is the mixed fraction of the second subpopulation, $\mathcal{PS}$ and $\mathcal{PL}$ are the PS and power-law primary-mass function, and $\mathcal{G}_{[-1,1]}$ is the truncated Gaussian distribution. $P(m_{2}|m_{1};{\bf\Lambda})$ is the normalized secondary-mass distribution conditioned on $m_{1}$, $P(\chi_{\rm eff}|q;\mu_{\chi,0},\sigma_{\chi,0},a,b)$ is the Base model that encodes the $\chi_{\rm eff}-q$ correlation (Callister et al., 2021), see Appendix B for the definition.
The other model has a more concise formula, i.e., the two $\chi_{\rm eff}$-distributions are modulated by a transition function of the primary mass (see also Wang et al., 2022; Li et al., 2024b; Guo et al., 2024, for similar constructions), hereafter Transition model,
$\pi_{\rm tran}$ $(m_{1},m_{2},\chi_{\rm eff}|{\bf\Lambda})=\mathcal{PS}(m_{1}|{\bf\Lambda})P(m_{2}|m_{1};{\bf\Lambda})$ (4)
$P_{\rm tran}(\chi_{\rm eff}|q,m_{1};{\bf\Lambda}),$
with
$P_{\rm tran}(\chi_{\rm eff}$ $|q,m_{1};{\bf\Lambda})=$ (5)
$P(\chi_{\rm eff}|q;\mu_{\chi,0},\sigma_{\chi,0},a,b)\frac{1}{1+e^{(m_{1}-m_{\rm t})/\delta_{\rm t}}}+$
$\mathcal{G}_{[-1,1]}(\chi_{\rm eff}|\mu_{\chi,2},\sigma_{\chi,2})\frac{1}{1+e^{(m_{\rm t}-m_{1})/\delta_{\rm t}}},$
where $m_{\rm t}$ and $\delta_{\rm t}$ are the location and rapidness of the transition. For the BBHs with primary mass below / above $m_{\rm t}$, the $\chi_{\rm eff}$ distribution is dominated by the Base model / second Gaussian distribution.
3 Results
Table 1: Model comparison
Model $\ln{\mathcal{B}}$
Base(PS) 0
Base(PS) with $a=0$,$b=0$ 0.8
Transition(PS) 3.1
Transition(PS) with $a=0$,$b=0$ 5.2
Mixture(PS) 2.1
Mixture(PS) with $a=0$,$b=0$ 4
Base(PP) -5.9
Base(PP) with $a=0$,$b=0$6.9
Transition(PP) with $a=0$,$b=0$3
Note: All the log Bayes factors are relative to the Base model
that encode the $\chi_{\rm eff}$-$q$ correlation and accompanied with PS
mass fucntion.
Figure 1: Constraints on the mean $\mu_{\chi}(q)$ and standard deviation $\sigma_{\chi}(q)$ of the $\chi_{\rm eff}$ distribution, as a function of BBH mass ratio $q$. The solid curves are the medians and the colored bands are the 90% credible intervals.
Figure 2: Posteriors of the hyperparameters describing the $\chi_{\rm eff}-q$ correlation. The contours mark the central 50% and 90% posterior credible regions, the values represent the median and 90% credible intervals.
In Table 1, we summarize the Bayes factors of the novel models in this work compared to the Base model. The Transition model and Mixture model are more favored than the Base model, with $\ln\mathcal{B}=3.1$ and $\ln\mathcal{B}=2.1$. Furthermore, the Transition model and Mixture model without the $\chi_{\rm eff}$-$q$ correlation (i.e., $a=0,b=0$) are even more favored. Figure 2 shows the posteriors of the hyperparameters describing the $\chi_{\rm eff}$ and $q$ correlation, the parameters of the Transition model and Mixture model highly consistent with each other. In both models, the slope parameters $a$ and $b$ are consistent with zero, indicating there is no evidence for a $\chi_{\rm eff}-q$ correlation in the first subpopulation or the low-mass range. In subsequent analysis, we adopt the results inferred with $a=0,b=0$, although the posterior distributions of parameters are broadly consistent with the results inferred with variable $a$ and $b$, see Appendix C.
We find that the second subpopulation shows a significantly different $\chi_{\rm eff}$ distribution compared to the first subpopulation, see Figure 3. The divergence point (of the Transition model) in the primary-mass function is $m_{\rm t}=49^{+13}_{-10}M_{\odot}$, with a transition scale of $6^{+4}_{-4}M_{\odot}$. As for the Mixture model, the maximum mass of the first subpopulation is $\sim 45M_{\odot}$, which is consistent with the lower edge of the Pair-instability Supernova BH mass gap (Farmer et al., 2019). The minimum mass of the second subpopulation is $\sim 30M_{\odot}$. These results are consistent with our previous analysis using spin magnitudes of BBHs (Wang et al., 2022; Li et al., 2024b), see Appendix C for the distributions of other related parameters. The $\chi_{\rm eff}$ distributions of both subpopulations are not symmetric with respect to zero, indicating the presence of formation channels other than star clusters.
We find that $\mu_{\chi,2}>0$ at 98.7% credible level, and a symmetric $\chi_{\rm eff}$ distribution with respect to zero for the second population (i.e.,$\mu_{\chi,2}=0$) is disfavored by a Bayes factor of $\ln\mathcal{B}=1.6$ ($\mathcal{B}\sim 5$), which indicates that the hierarchical mergers cannot be produced (solely) by star clusters. The second $\chi_{\rm eff}$ distribution is broad and peaks at $\mu_{\rm eff,2}=0.34^{+0.17}_{-0.22}$, which is typical for the hierarchical mergers in the AGN disks (Yang et al., 2019). We have also inferred with variable edges ($\chi_{\rm min}$, $\chi_{\rm max}$) truncated on the second $\chi_{\rm eff}$ distribution, see Appendix D. Though the case $\chi_{\rm min}=0$ cannot be ruled out, the location of the lower edge prefers a negative value. Therefore, it is likely that the star clusters also contribute to the hierarchical mergers.
Figure 3: Effective-spin distributions of the two subpopulation inferred by the Transition and Mixture models with $a=0$,$b=0$. The lines represent the mean values and 90% credible intervals
4 Conclusions and Discussion
In this work, we investigate the origins of $\chi_{\rm eff}$-$q$ correlation in the BBHs with data of GWTC-3. Different from the previous analysis (Callister et al., 2021; Abbott et al., 2023a), we adopt a more flexible primary-mass function (e.g., PS; Edelman et al., 2022). We find the evidence for the $\chi_{\rm eff}$-$q$ anti-correlation inferred with the PS model is weaker than that inferred with the PP model (Talbot & Thrane, 2018; Abbott et al., 2021b). This indicates that the parametric formula PP may introduce bias in the $\chi_{\rm eff}$-$q$ distribution, since there are additional structures in the primary-mass distribution beyond one power law plus one peak (Abbott et al., 2023a; Tiwari, 2022; Edelman et al., 2023; Callister & Farr, 2024).
With the dedicated models that introduce a second $\chi_{\rm eff}$ distribution for the second (high-spin) subpopulation, the inferred $\chi_{\rm eff}$-$q$ correlation vanishes. Therefore, we can conclude that the $\chi_{\rm eff}$-$q$ anti-correlation of the entire BBH population mainly results from the superposition of two subpopulations. Additionally, there is no evidence for a $\chi_{\rm eff}-q$ correlation in each subpopulation identified by the previous work (Li et al., 2024b, a), see Appendix E.
What are the second subpopulation?. We point to the hierarchical mergers in gas-rich environments, such as AGN disks (Yang et al., 2019). Our previous works (Wang et al., 2022; Li et al., 2024b) identified a subpopulation of hierarchical mergers in BBHs, providing smoking-gun evidence, specifically a spin-magnitude distribution of approximately 0.7 (Gerosa & Fishbach, 2021). Antonini et al. (2024) attributed this subpopulation to hierarchical mergers in star clusters, as they found that the $\chi_{\rm eff}$ distribution is consistent with a uniform distribution in the range of approximately (-0.5, 0.5). In this work, we employ a more flexible model and find that the high-spin subpopulation prefers an asymmetric $\chi_{\rm eff}$ distribution (with respect to zero) over a symmetric one, with a logarithmic Bayes factor of $\ln\mathcal{B}=1.6$ (where $\mathcal{B}\sim 5$). we also infer that the peak of the second $\chi_{\rm eff}$ distribution $\mu_{\chi,2}>0$ at 98.7% credible level.
Many simulations (e.g., Yang et al., 2019; Santini et al., 2023; Cook et al., 2024) show that the hierarchical mergers in disk-like environments can produce a $\chi_{\rm eff}$ distribution peaking at $\sim 0.35$, which is consistent with our results, $\mu_{\chi,2}=0.35^{+0.17}_{-0.22}(0.30^{+0.19}_{-0.19})$, from the Transition (Mixture) model. However, contribution of formation channels that can produce significantly negative $\chi_{\rm eff}$ values, such as hierarchical mergers in star clusters, cannot be ruled out, see Appendix D.
Kick velocities of the merger remnants are also helpful in determining the formation environments of hierarchical mergers (Gerosa & Fishbach, 2021; Zevin & Holz, 2022; Li & Fan, 2024). Several population analyses suggested that the excess of BBHs in $\sim 30-40M_{\odot}$ are consistent with the dynamical formation channels (Wang et al., 2022; Li et al., 2024a; Ray et al., 2024), and the ratio of merger rates between hierarchical mergers and first-generation (dynamical) BBHs are consistent with the star cluster origins (Antonini et al., 2023). Therefore, it is possible that the hierarchical mergers detected by LIGO/Virgo/KAGRA originate from both formation environments. However, there are only $\sim 11$ hierarchical mergers (high-spin events) in GWTC-3 (Li et al., 2024b), making it difficult to determine the mixture fractions of AGN-disk like and star-cluster formation channels. We will address such issue when the GW data are significantly enriched.
Does any single formation channel contributes to $\chi_{\rm eff}-q$ correlation? (Banerjee & Olejak, 2024) We find there is no evidence for a $\chi_{\rm eff}-q$ correlation in the low-mass sub-populations, which are associated with the first-generation BBHs (Li et al., 2024b). However, this sub-population may not originate from a single formation channel. Both isolated and dynamical formation channels have contributed to first-generation BBHs (Wang et al., 2022; Li et al., 2024b; Godfrey et al., 2023; Li et al., 2024a; Ray et al., 2024). Therefore, we further investigate the $\chi_{\rm eff}-q$ correlation of BBHs originating from potential formation channels (Wang et al., 2022; Li et al., 2024b; Godfrey et al., 2023; Li et al., 2024a; Ray et al., 2024) and find no evidence for either subpopulation. See Appendix E for details. Interestingly, we find a stronger $\chi_{\rm eff}-q$ correlation in the potential subpopulation associated with the dynamically formed BBHs (i.e., excluding the BBHs from isolated channels as indicated by Li et al. (2024a)). This indicates that the $\chi_{\rm eff}-q$ correlation originates from dynamical formation channels, and that the AGN-disk formation channels may take a significant contribution (Santini et al., 2023; Cook et al., 2024). The stable mass transfer formation channel is expected to produce BBHs with $\chi_{\rm eff}-q$ anti-correlation (Olejak et al., 2024; Banerjee & Olejak, 2024), while BBHs from the common envelop channel may exhibit an opposite correlation (Bavera et al., 2020, 2021). The $\chi_{\rm eff}-q$ correlation in the isolated channels may be measured when the detections are enriched and the sup-populations are determined.
The fourth observing run (O4) of the LIGO-Virgo-KAGRA GW detectors is currently underway, and the number of detections is rapidly increasing (see https://gracedb.ligo.org/latest/). At the end of O4, more than four times as many events are expected to be observed compared to O3 (Callister, 2024). With the enriched data more subpopulations / formation channels of BBHs may be identified (Zevin et al., 2021), and the mixture fractions and the parameter correlations of subpopulations can be better determined (Li et al., 2022, 2024a; Guo et al., 2024; Heinzel et al., 2024; Callister, 2024).
This work is supported by the National Natural Science Foundation of China (No. 12233011), the General Fund (No. 2024M753495) of the China Postdoctoral Science Foundation, and the Priority Research Program of the Chinese Academy of Sciences (No. XDB0550400). Y-Z Wang is supported by the National Natural Science Foundation of China (No. 12203101) and S-P Tang is supported by the National Natural Science Foundation of China (No. 12303056), the General Fund (No. 2023M733736) and the Postdoctoral Fellowship Program (GZB20230839) of the China Postdoctoral Science Foundation. This research has made use of data and software obtained from the Gravitational Wave Open Science Center (https://www.gw-openscience.org), a service of LIGO Laboratory, the LIGO Scientific Collaboration and the Virgo Collaboration. LIGO is funded by the U.S. National Science Foundation. Virgo is funded by the French Centre National de Recherche Scientifique (CNRS), the Italian Istituto Nazionale della Fisica Nucleare (INFN) and the Dutch Nikhef, with contributions by Polish and Hungarian institutes.
Appendix A Hierarchical Bayesian Inference
We perform hierarchical Bayesian inference to infer the hyperparameters $\Lambda$ describing population models $\pi(\theta|\Lambda)$. Following the framework described in (Mandel et al., 2019; Abbott et al., 2021b, 2023a), given $\Lambda$, the likelihood of the GW data $\{d\}$ from $N_{\rm det}$ detections can be expressed as,
[…]
Following Abbott et al. (2023a); Talbot & Golomb (2023), we define the effective number of samples for the $i$-th event in the Monte Carlo integral as $N_{{\rm eff},i}=\frac{[\sum_{j}{w_{i,j}]^{2}}}{\sum_{j}{w_{i,j}^{2}}}$, where $w_{i,j}$ is the weight of $j$-th sample in $i$-th event. We constrain $N_{{\rm eff},i}>10$ to ensure accurate evaluation of likelihood, which is sufficiently high given the sample size of GWTC-3 (Essick & Farr, 2022). Additionally, we constrain the effective number of found injections remaining after population reweighting as $N_{\rm eff,sel}>4N_{\rm det}$, to ensure an accurate estimation of $N_{\rm exp}$ (Farr, 2019; Abbott et al., 2021b).
Appendix B Population models
B.1 Mass function
The parametric primary-mass function is the popular PowerLaw+Peak model (Abbott et al., 2021b) , which reads,
$\mathcal{PP}$ $(m_{1}|{\bf\Lambda})\propto$ (B1)
$[\mathcal{PL}(m_{1}|-\alpha,m_{\rm min},m_{\rm max})(1-\lambda_{\rm peak})+$
$\mathcal{G}_{[m_{\rm min},m_{\rm max}]}(m_{1}|\mu_{\rm m},\sigma_{\rm m})\lambda_{\rm peak}]\times\mathcal{S}(m_{1}|\delta_{\rm m},m_{\rm min}),$
where $\mathcal{PL}(m_{1}|-\alpha,m_{\rm min},m_{\rm max})$ is the Power-law distribution with slope index of $-\alpha$ truncated on ($m_{\rm min}$, $m_{\rm max}$). $\mathcal{G}$ is the Gaussian distribution with mean $\mu_{\rm m}$ and width $\sigma_{\rm m}$ truncated on ($m_{\rm min}$, $m_{\rm max}$), $\mathcal{S}(m_{1}|\delta_{\rm m},m_{\rm min})$ is the smooth function with smooth scale of $\delta_{\rm m}$, impacting on the low edge $m_{\rm min}$.
In order to reduce the bias that may be brought by the mis-specification of parametric formulas, we use a semi-/non-parametric mass function(Edelman et al., 2022) for the main analysis,
$\mathcal{PS}$ $(m_{1}|{\bf\Lambda})\propto\mathcal{PL}(m_{1}|-\alpha,m_{\rm min},m_{\rm max})$ (B2)
×$\times\mathcal{S}(m_{1}|\delta_{\rm m},m_{\rm min})e^{f(m|\{x_{i}\},\{f_{i}\})},$
where $f(m|\{x_{i}\},\{f_{i}\})$ is the cubic-spline perturbation function interpolated between the knots ($x_{i},f_{i}$) placed in the mass range. Here we adopted 12 knots $\{x_{i}\}_{i=0}^{12}$ linearly distributed in log space of (6, 80)$M_{\odot}$, and restrict the perturbation to zero at the minimum and maximum knots.
The secondary-mass function is conditioned on the primary mass (Abbott et al., 2021b),
$P(m_{2}|m_{1};{\bf\Lambda})\propto\mathcal{PL}(m_{2}|\beta,m_{\rm min},m_{1})\mathcal{S}(m_{2}|\delta_{\rm m},m_{\rm min}).$ (B3)
B.2 Effective-spin distribution model
We follow Callister et al. (2021) to construct a mass-ratio-dependent model for effective-spin distribution, i.e., the Base model,
$P$ $(\chi_{\rm eff}|q;\mu_{\chi,0},\sigma_{\chi,0},a,b)=$ (B4)
$\mathcal{G}_{[-1,1]}(\mu_{\chi}(q;\mu_{\chi,0},a),\sigma_{\chi}(q;\sigma_{\chi,0},b)),$
with
$\mu_{\chi}(q;\mu_{\chi,0},a)=\mu_{\chi,0}+a(q-0.5),$ (B5)
$\log\sigma_{\chi}(q;\sigma_{\chi,0},b)=\log\sigma_{\chi,0}+b(q-0.5),$
where $\mathcal{G}_{[-1,1]}$ is the Gaussian truncated on $[-1,1]$. Note that the linear functions in Abbott et al. (2023a) is slightly different, where the authors let $\mu_{\chi}(q=1)=\mu_{\chi,0}$ and $\log\sigma_{\chi}(q=1)=\log\sigma_{\chi,0}$.
B.3 Rate evolution model
The merger rate density as a function of redshift reads (MD model Madau & Dickinson, 2014),
$R(z|\gamma,\kappa,z_{\rm p})=R_{0}\times\frac{[(1+z_{\rm p})^{(\gamma+\kappa)}+1](1+z)^{\gamma}}{(1+z)^{(\gamma+\kappa)}+(1+z_{\rm p})^{(\gamma+\kappa)}},$ (B6)
where $R_{0}$ is the local merger rate density. Note the injection campaign only provides mock events with $z<1.9$, so we normalize the redshift distribution $P(z|\gamma,\kappa,z_{\rm p})$ within $(0,1.9)$, when calculating likelihood.
Table 2: Summary of model parameters.
Parameter Description Prior
$m_{\rm min}[M_{\odot}]$ The minimum mass $U(2,10)$
$m_{\rm max}[M_{\odot}]$ The maximum mass $U(30,100)$
$\alpha$ Slope index of the power-law mass function $U(-8,8)$
$\delta_{\rm m}[M_{\odot}]$ Smooth scale of the mass lower edge $U(0,10)$
$\beta_{q}$ Slope index of the mass-ratio distribution $U(-8,8)$
Special for PowerLaw+Peak
$\lambda_{\rm peak}$ Fraction in the Gaussian component $U(0,1)$
$\mu_{\rm m}[M_{\odot}]$ Mean of the Gaussian component $U(20,50)$
$\sigma_{\rm m}[M_{\odot}]$ Width of the Gaussian component $U(1,10)$
Special for PowerLaw+Spline
$\{f_{i}\}_{i=2}^{11}$ Interpolation values of perturbation for mass function $\mathcal{N}(0,1)$
Effective-spin distribution model
$\mu_{\chi,0}$ Mean of $\chi_{\rm eff}$ distribution given $q=0.5$ $U(-1,1)$
$\lg\sigma_{\chi,0}$ Log width of $\chi_{\rm eff}$ distribution given $q=0.5$ $U(-1.5,0.5)$
$a$ Correlation between $\mu_{\chi}$ and $q$ $U(-2.5,2.5)$
$b$ Correlation between $\log\sigma_{\chi}$ and $q$ $U(-2,2)$
Rate evolution model
$\lg(R_{0}[{\rm Gpc}^{-3}~{}{\rm yr}^{-1}])$ Local merger rate density $U(-3,3)$
$z_{p}$ Peak point for the rate evolution function $U(0,4)$
$\gamma$ Slope of the power-law regime before $z_{p}$ $U(-8,8)$
$\kappa$ Slope of the power-law regime after $z_{p}$ $U(-8,8)$
Transition model
$m_{\rm t}[M_{\odot}]$ The transition point in primary-mass function $U(20,70)$
$\delta_{\rm t}[M_{\odot}]$ Smooth scale of the transition $U(0,10)$
$\mu_{\chi,2}$ Mean of $\chi_{\rm eff}$ distribution for secondary subpopulation $U(-1,1)$
$\lg\sigma_{\chi,2}$ Log width of $\chi_{\rm eff}$ distribution secondary subpopulation $U(-1.5,0.5)$
Mixture model
$m_{\rm min,2}[M_{\odot}]$ Minimum mass of the secondary component $U(10,50)$
$m_{\rm max,2}[M_{\odot}]$ Maximum mass of the secondary component $U(50,100)$
$\alpha_{2}$ Index of $m_{1}$ distribution in the secondary component $U(-8,8)$
$\mu_{\chi,2}$ Mean of $\chi_{\rm eff}$ distribution for secondary subpopulation $U(-1,1)$
$\lg\sigma_{\chi,2}$ Log width of $\chi_{\rm eff}$ distribution secondary subpopulation $U(-1.5,0.5)$
$r_{2}$ fraction of the secondary component $U(0,1)$
Note: $U$ and $\mathcal{N}$ are for Uniform and Normal distribution.
Appendix C Additional results
Figure 4 and Figure 5 present the full hyperparameters (except for the interpolation knots) of the Transition model, and Mixture model respectively. We find that whether we fix $a=0,b=0$ or not does not affect the transition / classification of the two subpopulations. These results are consistent with the analysis using spin-magnitude distributions in our previous work (Li et al., 2024b). We observe that the results of the two models are well consistent with each other. The first $\chi_{\rm eff}$ distribution narrowly peaks at $\sim 0.05$, while the second $\chi_{\rm eff}$ distribution peaks at $\sim 0.35$, favoring a contribution from hierarchical mergers in AGN disks (Yang et al., 2019).
Figure 4: Posteriors of the special hyperparameters for the Transition model. The contours mark the central 50% and 90% posterior credible regions, the values represent the median and 90% credible intervals.
Figure 5: Posteriors of the special hyperparameters for the Mixture model. The contours mark the central 50% and 90% posterior credible regions, the values represent the median and 90% credible intervals.
Appendix D Is the second $\chi_{\rm eff}$ distribution symmetric?
The symmetry of $\chi_{\rm eff}$ distribution is critical for determining the formation environments of hierarchical mergers. Hierarchical mergers in star clusters always exhibit a symmetric $\chi_{\rm eff}$ distribution (Payne et al., 2024; Fishbach et al., 2022). While AGN-driven hierarchical mergers tend to favor positive $\chi_{\rm eff}$ due to gas torques (McKernan et al., 2018). In this section, we infer with variable lower edge and upper edge ($\chi_{\rm min,2}$, $\chi_{\rm max,2}$) for the second $\chi_{\rm eff}$ distribution. We infer that $\chi_{\rm max,2}>0.4$ at 97% credible level. Although, we cannot rule out the case of $\chi_{\rm min,2}=0$, it is more likely that $\chi_{\rm min,2}$ takes on a negative value, as shown in Figure 6.
Figure 6: $\chi_{\rm eff}$ distributions inferred using Transition with variable edges $\chi_{\rm min,2}$ and $\chi_{\rm max,2}$ truncated on the second $\chi_{\rm eff}$ distribution.
Appendix E Are there $\chi_{\rm eff}-q$ correlation in subpopulations?
As presented in Section 3, we find no evidence for $\chi_{\rm eff}-q$ correlation in the low-spin (or first-generation) subpopulation. However, the first-generation subpopulation is most likely a mix of BBHs originating from both isolated and dynamical channels (Wang et al., 2022; Li et al., 2024a; Ray et al., 2024). Therefore, it is possible that the $\chi_{\rm eff}-q$ correlation occurs in a single, specific channel, such as the isolated evolution channel (Banerjee & Olejak, 2024).
We first artificially divide the entire population into three subpopulations according to our previous works (see Supplemental Table IV of Li et al., 2024b) and (Table 5 of Li et al., 2024a). The first is the low-spin & nearly-aligned subpopulation, that is related to the isolated evolution channels, see Table 4. The second is the high-spin subpopulation, related to hierarchical mergers, which are also identified within this work. The third is the low-spin & isotropic subpopulation that is associated with the first-generation BBHs formed in star clusters. The reason why we do not jointly fit the three subpopulation in this work is that, it is difficult to distinguish the first and third subpopulations using the $\chi_{\rm eff}$ distribution, since the effective spins in both subpopulations are close to zero as shown in Figure 3.
We then fit the three sets of samples with the Base (PS) model, respectively. Figure 7 shows the posterior distributions of the parameters describing the $\chi_{\rm eff}$ and $q$ distribution in each potential subpopulation. We find there is no evidence of correlation in either subpopulation, see also Table 3.
Interestingly, if we fit the combination of the second and third sub-set samples, which is potentially associated with the dynamical assembly (Li et al., 2024a), then we find strong evidence for the $\chi_{\rm eff}$-$q$ correlation with $\ln\mathcal{B}=2.6$, see also Figure 7 (Right). This result is consistent with the predictions of mergers in AGN disks (Cook et al., 2024). We find that a significant fraction of BBHs in the 2nd pop have large (positive) $\chi_{\rm eff}$ values, while the BBHs in the 3rd pop are located at $\chi_{\rm eff}\sim 0$,$q\sim 1$. Therefore, the $\chi_{\rm eff}$-$q$ anti-correlation likely arises from the superposition of these two types of subpopulations.
Figure 7: Left: posterior distribution of parameters describing the $\chi_{\rm eff}$-$q$ distribution in each subpopulation. Right: posterior $\chi_{\rm eff}$-$q$ distribution of individual event reweighed to the population-informed prior. The shaded areas mark the 90% credible regions and the points stand for the mean values.
Table 3: Model comparison
Data Sets $\ln{\mathcal{B}^{\rm with~{}corr.}_{\rm w./o.~{}corr.}}$
Whole -0.8
Low-Spin Aligned (Field; 1st pop) -1.2
High-Spin (2G; 2nd pop) -0.1
Low-Spin Isotropic (Dynamical 1G; 3rd pop) -1.0
Low-Spin (1G; 1st+3rd pop) -2.1
Low-Spin Isotropic + High-Spin (Dynamical; 2nd+3rd pop) 2.6
Note: Bayes factor for each potential subpopulation with versus without a $\chi_{\rm eff}-q$ correlation.
Table 4: Events of each potential subpopulation
Low-Spin Aligned (Field; 1st pop)
GW151012_095443 GW151226_033853 GW170608_020116 GW190412_053044 GW190512_180714 GW190707_093326
GW190708_232457 GW190720_000836 GW190728_064510 GW190828_065509 GW190924_021846 GW190930_133541
GW190725_174728 GW191105_143521 GW191129_134029 GW191204_171526 GW191216_213338 GW200202_154313
GW200316_215756 GW191103_012549
High-Spin (2G; 2nd pop)
GW170729_185629 GW190517_055101 GW190519_153544 GW190521_030229 GW190602_175927 GW190620_030421
GW190701_203306 GW190706_222641 GW190929_012149 GW190805_211137 GW191109_010717
Low-Spin Isotropic (Dynamical 1G; 3rd pop)
GW150914_095045 GW170104_101158 GW170809_082821 GW170814_103043 GW170818_022509 GW170823_131358
GW190408_181802 GW190413_134308 GW190421_213856 GW190503_185404 GW190513_205428 GW190521_074359
GW190527_092055 GW190630_185205 GW190727_060333 GW190803_022701 GW190828_063405 GW190910_112807
GW190915_235702 GW190925_232845 GW190413_052954 GW190719_215514 GW190731_140936 GW191127_050227
GW191215_223052 GW191222_033537 GW191230_180458 GW200112_155838 GW200128_022011 GW200129_065458
GW200208_130117 GW200209_085452 GW200219_094415 GW200224_222234 GW200225_060421 GW200302_015811
GW200311_115853 GW200216_220804
Note: The 2nd sub-set of GW events are figured out according to Li et al. (2024b), each event has probability $>0.5$ to
containing at least one high-spin BH. The 3rd sub-set are figured out according to Li et al. (2024a), each event has
probability $>0.5$ belonging to isotropic-spin sub-population.